OpenStudy (anonymous):

Can you check these please? Thank you! Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = -3/5; y-intercept = (0, 2) slope = 5/3; y-intercept = (0, -6) I have know idea about this one I think it's B but for some reason this one really confuses me. Use the distributive property to simplify the expression below: -4(3u - 5) -12u - 5 -12u + 20 -12u - 20 -12u + 9 I think it's B Solve: 6x - 2 > 8x + 4 x > -3 x -3 x < -3 x -3 I think C.

OpenStudy (rajee_sam):

problem no. 2 and 3 are correct. Problem No. 1 is not

OpenStudy (rajee_sam):

for problem no.1 you are given the equation of a line in the standard form which is Ax + By = C This you have to convert to Slope-Intercept Form which is y = mx + b, where m = Slope and b = y-intercept. Can you do that? Change the standard form to slope-intercept form.

OpenStudy (anonymous):

I don't know I'm horrible at slopes could you help me a bit more?

OpenStudy (rajee_sam):

we are not doing anything with slopes per se. All you are doing is solving the given equation for y. You are given an equation 5x -3y = -6 This equation you have to solve for y. meaning you have to isolate the y and rewrite the whole equation as y = "something" . Can you do that? Once you do that we can identify the slope and y-intercept from the resulting equation.

OpenStudy (anonymous):

Okay thank you.

OpenStudy (anonymous):

Would it be A?

OpenStudy (texaschic101):

for #1, you have to put it in " y = mx + b " form and m would be the slope and b would be the y-intercept. Lets do that..... 5x - 3y = -6 (you want -3y on one side and everything else on the other side) -3y = -5x - 6 (now you want to divide by -3 to get y by itself) y = 5/3x + 2 slope being 5/3 and y-intercept being 2 -4(3u - 5) you distribute the -4 through -4 * 3y + -4 * -5 = -12y + 20 6x - 2 > 8x + 4 you want to get the x's on one side, so first add 2 to both sides 6x - 2 + 2 > 8x + 4 + 2 (combine like terms 6x > 8x + 6 (now you want to subtract 8x from both sides) 6x - 8x > 8x - 8x + 6 (combine like terms) -2x > 6 (now you have to divide by -2, keeping in mind when you divide by a negative, the inequality sign changes) -2x/-2 < 6/-2 x < -3 any questions ?